GlueX Start Counter Meeting, April 28, 2015
GlueX Start Counter Meeting
Tuesday, April 28, 2015
10:00 am EDT
CEBAF Center L102
- Collaboration Meeting
- Review minutes from the April 16 meeting
- Detector Status
- DAQ Tasks for Spring 2015 Commissioning Run
- Meeting Schedule
- Action Item Review
- The BlueJeans Meeting ID is 556 286 544 .
- Connect to the Meeting
Talks can be deposited in the directory
/group/halld/www/halldweb/html/talks/2015 on the JLab CUE. This directory is accessible from the web at https://halldweb.jlab.org/talks/2015/ .
- FIU: Werner Boeglin, Mahmoud Kamel
- JLab: Mark Ito (chair), Eric Pooser, Beni Zihlmann
News from the Run
- With a 200 nA beam on a 3×10-4 radiator, Eric reported see rates from 100 to 200 kHz in the individual start counter sectors.
- When beam started up last Friday, the scalers rate distribution was symmetric, but recent log entries show an asymmetry of 10 to 20% in the rates. The cause is not understood at present. It is nothing like the asymmetry we saw last Fall when the start counter was tilted by about 6 degrees. Werner remarked that we should expect some level of asymmetry in the raw scalers rates.
- Werner asked about whether thresholds in the discriminators can be set on a channel-by-channel basis. Eric told us that the boards accept different thresholds, but our current software can only load a single value for all channels on a board.
- Bias voltages on the SiPMs are currently all set as per Hamamatsu's recommendation, as tabulated by Nick and Fernando. Recall the four SiPMs for a given sector are all run from a common voltage channel.
Bias Voltage Scan
We discussed a plan to scan bias voltages for the SiPMs. This has never been done for the start counters, including during fabrication and testing at FIU. The study would tell us if we should pursue gain balancing among the sectors. Eric reminded us that the gain should change by about one percent per 10 mV change in bias voltage. We decided on a plan to vary the voltages in the range of -200 mV to +200 mV from nominal over the next few days, changing setting after we accumulate a significant chunk of data (a few hours of beam). Thus the gain should only change in the range of -20% to 20% during this exercise.
Calibration work is still in progress, both the time-walk correction and propagation delay correction scheme are evolving. Eric also mentioned that there will be some work to incorporate the new calibration schemes in the PID library.
We agreed to go to a two-week schedule, starting after the collaboration meeting.